Solve ∫y=0^{1}∫x=y^{2}^{1}∫z=0^{1-x}x dzdxdy


1
Expert's answer
2021-08-11T07:28:33-0400

"\int _0^1\int _{b^2}^1\int _0^{1-x}xdzdxdy\\
=\int _0^1\int _{y^2}^1x\left(-x+1\right)dxdy\\
=\int _0^1\left(\frac{1-y^4}{2}-\frac{1-y^6}{3}\right)dy\\
=\frac{4}{35}"